# Mergers Fall Short: Non-merger Channels Required for Galactic Heavy Element Production

### v2

@@ IV Discussion and Conclusion
We have investigated whether fast-merging BNSs or NSBH mergers can resolve the -process enrichment discrepancy identified previously Côté (2019); Cowan et al. (2021); Kobayashi et al. (2023); van de Voort et al. (2015); Siegel et al. (2019b); Chen et al. (2025), where BNS mergers fail to reproduce the observed Eu trends—the negative slope in [Eu/Fe] versus [Fe/H]—in the Milky Way disk. Across all three scenarios explored—GW-detectable fast-merging BNSs, GW-undetectable fast-merging BNSs, and NSBH mergers—we find that these channels are unlikely to serve as the dominant source of -process enrichment.
@@
We emphasize that, although the fast-merging BNS and NSBH populations may not precisely follow the assumed star formation rate, may have different yields from those estimated with semi-analytical fit, may have mass distributions that differ from those inferred from GW observations, and/or may not adhere to the assumed -process abundance pattern—all of which could affect their event rates or Eu yields—our analysis still indicates that an exceptionally large contribution from these channels is required across cosmic history (see the required fractions and volumetric rates in the Results section). This poses a substantial challenge to constructing a consistent theoretical framework. For instance, to reduce the event rates of fast-merging BNSs or NSBHs to a level comparable to that of delayed BNSs, their Eu yield would need to be increased by a factor of . This requires a scenario that boosts their yields by times without increasing the delayed BNS yield. Similarly, even if the early-Universe evolution of these populations differs from our assumptions, they would still need to occur much more frequently than delayed BNSs to reproduce the observed trends in stellar metallicity.
@@ Acknowledgements.
In this appendix, we present additional plots that shed light on further details of each of the three analyses discussed in the main text. Figure 4 shows corner plots of the posterior distributions for all model parameters: the total BNS rate (including both delayed and fast–merging channels); the ejecta mass (assumed common to both channels); the delay–time parameters and of the delayed BNS channel; and the fraction , which specifies the portion of assigned to the fast–merging channel required to reproduce the observed abundance trends of Milky Way disk stars Battistini and Bensby (2016). The large panel in the upper right shows the 90% credible regions (orange) in the – plane implied by the posterior samples; observed abundances from Battistini and Bensby (2016) are overplotted as blue points.

### v3

@@ IV Discussion and Conclusion
We have investigated whether fast-merging BNSs or NSBH mergers can resolve the -process enrichment discrepancy identified previously Côté (2019); Cowan et al. (2021); Kobayashi et al. (2023); van de Voort et al. (2015); Siegel et al. (2019b); Chen et al. (2025), where BNS mergers fail to reproduce the observed Eu trends—the negative slope in [Eu/Fe] versus [Fe/H]—in the Milky Way disk. Across all three scenarios explored—GW-detectable fast-merging BNSs, GW-undetectable fast-merging BNSs, and NSBH mergers—we find that, under current theoretical and observational expectations for compact-binary populations, these channels are unlikely to serve as the dominant source of -process enrichment. This conclusion reflects the tension between the event rates required by our chemical-evolution inference and existing expectations for fast-merging and NSBH populations.
@@
We emphasize that, although the fast-merging BNS and NSBH populations may not precisely follow the assumed star formation rate, may have different yields from those estimated with semi-analytical fit, may have mass distributions that differ from those inferred from GW observations, and/or may not adhere to the assumed -process abundance pattern—all of which could affect their event rates or -process yields—our analysis still indicates that an exceptionally large contribution from these channels is required across cosmic history (see the required fractions and volumetric rates in the Results section). This poses a substantial challenge to constructing a consistent theoretical framework. For instance, to reduce the event rates of fast-merging BNSs or NSBHs to a level comparable to that of delayed BNSs, their yield would need to be increased by more than an order of magnitude. This requires a scenario that boosts their yields by an order of magnitude without increasing the delayed BNS yield. Similarly, even if the early-Universe evolution of these populations differs from our assumptions, they would still need to occur much more frequently than delayed BNSs to reproduce the observed trends in stellar metallicity. To explore a very different SFR, we repeated our analysis assuming a constant SFR and found that our conclusions remain unchanged (see Appendix-B and Fig. 7 for more details).
While we have assumed identical r-process yields for delayed and fast-merging BNSs, this assumption is not required. In a more general scenario, the yields from these two channels may differ. In particular, systems classified as fast-merging may preferentially form on highly eccentric orbits, perhaps leading to more violent encounters than quasi-circular binaries and potentially producing a larger mass (25-30% more) of neutron-rich ejecta, as suggested by previous studies Papenfort et al. (2018); Foucart et al. (2024). To explore this possibility, we repeat our fast-merging BNS analyses assuming a yield that is twice that of delayed BNSs (100% more). Even under this optimistic assumption, our results indicate that fast-merging BNSs must still dominate the total merger rate: they must contribute more than 94% of the total rate in the scenario where GW searches are sensitive to fast mergers, and require a local-universe rate enhancement of when GW searches are not sensitive to fast mergers.
We note that this conclusion is also supported by independent astrophysical considerations regarding the environments in which fast-merging BNS systems are expected to form. Such systems are commonly associated with dense stellar environments, including globular clusters and nuclear star clusters. In the Milky Way, globular clusters are predominantly old, metal-poor, and reside in the halo, with star formation ceasing well before the formation of the thin disk considered here, making efficient enrichment of disk stars unlikely. Nuclear star clusters, while more metal-rich, are spatially concentrated near the Galactic center and are similarly disfavored as major contributors to the r-process content of the solar neighborhood. These considerations provide complementary physical support for our conclusion that fast-merging BNSs are unlikely to dominate r-process enrichment in the Galactic thin disk.
@@ Acknowledgements.
In this appendix, we present additional plots that shed light on further details of each of the three analyses discussed in the main text. Figure 4 shows corner plots of the posterior distributions for all model parameters: the total BNS rate (including both delayed and fast–merging channels); the ejecta mass (i.e., the per-event -process yield, assumed common to both channels); the delay–time parameters and of the delayed BNS channel; and the fraction , which specifies the portion of assigned to the fast–merging channel required to reproduce the observed abundance trends of Milky Way disk stars Battistini and Bensby (2016). The large panel in the upper right shows the 90% credible regions (orange) in the – plane implied by the posterior samples; observed abundances from Battistini and Bensby (2016) are overplotted as blue points.
@@
Appendix B Impact of star formation history models
In this section, we assess the sensitivity of our results to the assumed SFR model. We repeat the full inference under a constant SFR—an extreme limiting case—and compare it to the cosmic SFR used in the main text. The three panels in Fig. 7 summarize the outcome for each secondary–channel scenario considered in the paper: (i) fast–merging BNS assumed detectable by current GW searches; (ii) fast–merging BNS assumed missed by current searches; and (iii) NSBH mergers. Across all three cases the posterior shifts between the two SFR assumptions are modest, and the qualitative conclusions of the paper are unchanged: neither fast–merging BNS nor NSBH mergers can account for the Milky Way’s -process enrichment. Bayesian model selection decisively prefers the cosmic SFR in every scenario, with logarithmic Bayes factors , , and , respectively, for the three cases. This justifies adopting the cosmic SFR in the main analysis and demonstrates that our main findings are robust to the SFR choice.